OpenStudy (anonymous):

Determine the elements in the domain and range of G = { ( x, y) | y=(x-2)^2+x and x is a positive integer less than 6 }. The range of G is a. {1,2,3,4,5} b. {2,4,8,14} c. {2,4,8,14,22} d. {1,2,3,4,5,6}

OpenStudy (zepp):



OpenStudy (zepp):

This is a parabola, you'll need to find what value are in the range you described

OpenStudy (zepp):

So, for the first value of the range (Range is the y-values), you simply plug in the 6 y=(6-2)^2+6 y=4^2+6 y=16+6 y=22

OpenStudy (zepp):